# The Design Space of Lockfiles Across Package Managers

### v1

@@ 3.5. Results for RQ2: To what extent do open-source projects include lockfiles in version control?
For Go, most projects (92%) begin including go.sum within six months of project creation. The only exceptions are 4 projects: GoogleCloudPlatform/buildpacks, eolinker/apinto,
go-admin-team/go-admin, and shalb/cluster.dev. These projects have 13, 48, 29, and 35 direct dependencies in their main modules, respectively. The developers of the project, go-admin-team/
go-admin
committed the lockfile at the early stages of the project but later decided to remove it from version control. The reason for removal is not documented in the public repository. The projects GoogleCloudPlatform/buildpacks and eolinker/apinto have go.sum files for some modules but do not contain one for the main module.
GoogleCloudPlatform/buildpacks also uses the Go vendor folder which allows committing dependencies alongside the project, making them accessible at anytime. Vendoring loads dependencies from the local vendor folder instead of resolving them from remote repositories, thus supporting air-gapped builds. Yet, the go.sum file would still be useful to ensure the integrity of the dependencies properly.
@@ 1. RQ3:
    What benefits do developers perceive in using lockfiles? Here, we study the benefits of lockfiles in terms of functionality, readability, and security. Moreover, we assess how developers’ perceptions of lockfiles vary across different package managers. Understanding the practical benefits of lockfiles helps clarify their role in secure and reliable software development.
@@ 2. RQ4:
    What challenges do developers face when managing lockfiles? With this RQ, we study the effort required to learn about lockfile related functionalities, the maintenance burden, and situations where lockfiles fall short of meeting developer expectations. These findings are beneficial for package manager maintainers to improve their lockfile implementations for better usability and effectiveness.
4.2. Methodology for RQ3 and RQ4
To answer our 2 research questions on lockfile usage, we select a set of projects from each ecosystem, including some projects that publish their lockfile in the version control system and some projects that do not publish it. We then conduct semi-structured interviews with one of the top contributors of each project. In the following, we describe the methodology we follow to select the projects, contact developers, interview them and analyze their responses.
@@ 4.2. Methodology for RQ3 and RQ4
For each selected project, we gather publicly available contact information for the top contributor and send an interview invitation. If contact information is unavailable, or we do not receive a response within a week, we continue iterating through our randomly selected projects, sending invitations until we reach our target of two or four participants per category, for a total of 14 responses. This approach allows us to select a diverse group of developers from various backgrounds and location, who have substantial expertise in dependency management and who contribute to projects in various application domains.
@@
Interviews: We conduct 20 minute online interviews with 15 selected developers. We record the interviews upon consent from the participants. One interview (P8) was conducted in person at the participant’s request and lasted approximately an hour; in this case, only written notes were taken. The participants were not provided any compensation for their participation.
@@
Interview response analysis: We use the Whisper speech recognition tool (Radford et al., 2023) to transcribe 15 interviews with available audio recordings. If audio recordings are not available, we refer to the written notes. Two authors carry out line-by-line coding for the transcribed interviews (or written notes). Consequently, the same two authors collaboratively create an affinity diagram, grouping codes into emerging themes inductively. The final affinity diagram consists of 20 sub-themes and 7 main themes related to general software development practices, lockfile benefits, developer expectations, and challenges.
Experience (Years) Package Manager Lockfile Usage Project Stats Project Domain
@@
P1 ¿ 20 npm pnpm Yes 3k 110,892 320 Website development
@@
P6 2 - 5 Cargo Yes 1.9k 1338 50 Untracked social media
@@
P10 5 - 10 Poetry Pipenv Yes 834 512 27 Home Assistant Integration
@@
Table 7. Demographics of developers interviewed about their usage of lockfiles. The project domains are selected based on GitHub project descriptions. The experience indicates the number of years of experience of participants in programming in any language.
4.3. Results for RQ3: What benefits do developers perceive in using lockfiles?
@@
4.3.1. Build determinism benefit
@@ 4.3.2. Integrity benefit
Identifying package tampering through integrity verification requires that releases remain immutable. In Go, downloading packages directly from repositories like GitHub, where releases can be changed after publication, can lead to checksum failures if a package is updated under the same version. A Go developer described this as a trade-off between the convenience of releasing and the security. (”it’s good on the one hand that you will always, like from a security perspective, make sure that you will always get the code that is basically stored on the proxy. And if you don’t have that, something is wrong and the checksum will give you an error. But then again, it can cause issues if you like want to re-release the same version. And yeah, it’s like convenience versus security”, P15)
4.3.3. Transparency benefit
@@ 4.3.5. Security benefit
4.4. Results for RQ4: What challenges do developers face when managing lockfiles?
@@
4.4.1. Indirect dependency locking challenge
@@
4.4.2. Slowing down dependency updates
@@
4.4.3. Human readability challenge
@@
When asked about the information they consider during code reviews of lockfiles, the majority of developers who commit lockfiles mention that they focus on the name, URL, version of the package, and the number of indirect dependencies it includes. A consistent observation we make is that none of the developers pay attention to the checksum of the dependencies in the lockfile. (I’ve never actually checked checksums, P7). Following the same reasoning, Go developers recommend the Go’s approach of having checksums in a separate file. (”To be honest, I really don’t care where the checksum is stored. It makes it clean to have it separated because I only have to put it into the Go mod file and then the checksum will be generated automatically in the some file. So that’s like having a look at your Go mod file is much cleaner., P15 ”).
4.4.4. Cache challenge
@@ 4.4.5. Learning challenge
Findings from RQ4: Developers who do not publish their lockfile in version control, do so for clear reasons. These reasons include avoiding lockfiles when developing libraries, readability challenges due to the verbosity of lockfiles, or avoiding delays in dependency updates. Overall, developers agree on the fact that lockfile challenges are better managed in Go than in JavaScript or Python.
5. Informing the Future of Package Managers
@@
[Recommendation #1] We recommend committing lockfiles for all projects.
Commiting lockfiles allow for distributed integrity checks and deterministic builds. This holds for libraries and applications, as library developers face the same security risks and challenges of collaborative development as application developers. To avoid unexpected issues with in-range breakages that downstream users may face, application developers should be able to perform dependency resolution in different ways, with or without enforcing locked versions.
[Recommendation #2] We recommend future package managers to prioritize developer experience.
@@
[Recommendation #3] We recommend lockfiles to include only essential content.
When designing lockfiles, package managers should avoid duplicating information already present in the dependency specification file. They should also exclude data that is irrelevant to dependency locking. In npm, for example, including data such as licenses and funding information in the lockfile adds little to no value while unnecessarily increasing code reviewing cost. At the other extreme, Gradle presents a serious lockfile anti-pattern by omitting essential details such as checksums. We recommend lockfiles to include the following mandatory details: resolved package versions, checksums, URLs for downloaded packages, and a clear distinction between direct and indirect dependencies. Any additional metadata should be kept as brief as possible, and ideally included only for reproducibility purposes, such as OS, engine, and language versions. We also encourage separating checksums into a separate file, similar to the approach of Go, to improve the human readability of lockfiles.
[Recommendation #4] We command package managers to generate lockfiles by default.
@@
6. Related work
To the best of our knowledge, lockfiles have not been studied in detail in the literature, a surprising fact given their importance in the field. Some studies occasionally mention them in the broader context of dependency management.
6.1. Dependency Resolution and Package Managers
One of the key responsibilities of a package manager is resolving compatible versions for all the dependencies in a project. Zacchiroli et al. (Abate et al., 2020) review the usage of dependency resolving algorithms, including specialized constraint solving algorithms as well as ad-hoc dependency tree based algorithms. They conduct a census of dependency resolution mechanisms across 18 package managers, using version locking as one of the criteria for comparison. Pinckney et al. introduce a unified framework built on top of Max-SMT solvers to resolve dependencies more systematically, moving beyond ad-hoc algorithms (Pinckney et al., 2023).
Several studies focus on dependency conflict issues in different ecosystems. For example, (Wang et al., 2020) identifies factors that lead to conflicts, whereas (Wang et al., 2023) highlights issues related to inefficiency and excessive resource usage by the dependency resolution strategies in Python. (Wang et al., 2019) introduces a tool to generate crash stack traces for dependency conflicts in Java. Patra et al. (Patra et al., 2018) analyze dependency conflicts in JavaScript that arise from namespace collisions. (Wang et al., 2025) presents an in-depth study of conflicts between peer dependencies in npm. Similarly, in the Go ecosystem, Wang et al. (Wang et al., 2021) study conflicts caused by the coexistence of two library referencing modes: Go PATHs and Go modules.
Liu et al. (Liu et al., 2022) analyze vulnerability propagation within dependency trees by applying npm-specific dependency resolution rules, rather than relying on simple reachability analysis for indirect dependencies. They recommend managing dependencies with lockfiles. Along similar lines, He et al. (He et al., 2025) study the effect of dependency pinning in npm projects. They find that local pinning leads to more security vulnerabilities due to bloated and outdated dependencies. They also suggest that the risk of malicious package updates can be reduced when a small set of core dependencies pin their versions and keep them updated regularly.
@@ 6.1. Dependency Resolution and Package Managers
6.2. Dependency Management Practices in Open Source Software
Several empirical studies provide insights into the dependency management practices of open source software projects. Bogart et al. (Bogart et al., 2021) discuss practices related to breaking changes across different package managers, and highlight the use of lockfiles to lock versions of indirect dependencies as a means to stabilize projects. Breaking updates have been well studied in the literature (Ochoa et al., 2022; Reyes et al., 2024; Frank et al., 2024). Venturini et al. (Venturini et al., 2023) quantitatively evaluate the impact of breaking updates on dependent packages in the npm ecosystem. They also note the advantages of using lockfiles to track all resolved packages and avoid in-range breaking changes.
Schueller et al. (Schueller et al., 2022) curate data from the Rust ecosystem over eight years, capturing developer activity, library dependencies, and usage trends. Another empirical analysis in Rust (Li et al., 2023) reveals that 46% of packages adopted yanked releases and the proportion of yanked releases keeps increasing over the years. In Cargo, such yanked releases can be resolved only if a lockfile is present.
There is a body of work that analyzes security practices and threats associated with dependency management. Gu et al. (Gu et al., 2023) conduct a systematic analysis of security threats in package registries; Maven, PyPI, npm, Cargo, NuGet, and Go. Rahman et al. (Rahman et al., 2025) propose a metric to evaluate vulnerabilities in dependencies based on the time it takes for a fix to be applied. Bos (Bos, 2023) reviews supply chain attacks and discusses a type of attack that originates from lockfile tamperings. One such tampering is adding a duplicate entry to the lockfile with a modified package and its new checksum. As most lockfiles are tediously long for thorough review, such tampering goes undetected leading to supply chain attacks. Vaidya et al. (Vaidya et al., 2019) characterize attacks in npm and PyPI ecosystems. Their findings suggest that with proper tooling and practices it is possible to mitigate these attacks to a great extent. The authors discuss the npm CLI lockfile npm-shrinkwrap that allows indirect dependency locking. They mention that npm-shrinkwrap can prevent future attacks from new packages through dependency locking. Kabir et al. (Kabir et al., 2022) study security best practices in software development and consider committing lockfiles as one of the best practices.
@@ 6.2. Dependency Management Practices in Open Source Software
6.3. Adoption of Dependency Management Tools within Development Teams
Previous works have also used combined qualitative and quantitative methods to learn about developers’ adoption of dependency management tools. Schorlemmer et al. (Schorlemmer et al., 2024) study the adoption of software signing in four package registries; Maven, PyPI, DockerHub and Huggingface. Kalu et al. (Kalu et al., 2024) conduct a follow-up interview study to understand software signing in practice. They identify that strict signature rules increase the quantity of signatures, and that registry policies have an impact on the developer decision to adopt proper signing.
Saroar et al. (Saroar and Nayebi, 2023) survey 90 developers to understand how they use, choose, and troubleshoot GitHub Actions. Holtgrave et al. (Holtgrave et al., 2025) analyze different attack types that can be performed by impersonating GitHub user accounts. Bifolco et al. (Bifolco et al., 2024) assess the accuracy of the GitHub dependency graph in Java and Python projects. Bifolco et al. use lockfiles as one source of truth to assess the GitHub dependency graph.
Dependency management bots such as dependabot, renovate and Greenkeeper automate the process of updating dependency specification files and lockfiles when new dependency versions are available. Alfadel et al. (Alfadel et al., 2021) and Mohayeji et al. (Mohayeji et al., 2023) evaluate how developers respond to security updates suggested by dependabot. Rombaut et al. (Rombaut et al., 2023) study whether Greenkeeper reduces developer effort or introduces unnecessary workload. They mention lockfiles as a way to overcome in-range breaking changes, and also briefly mention about the challenge library developers face for sharing lockfiles. Our qualitative analysis confirms that developers do use bots to maintain dependency specification files and lockfiles up to date.

### v2

@@ 3.5 Results for RQ2: To what extent do open-source projects include lockfiles in version control?
For Go, most projects (92%) begin including go.sum within six months of project creation. The only exceptions are 4 projects: GoogleCloudPlatform/buildpacks, eolinker/apinto, go-admin-team/go-admin, and shalb/cluster.dev. These projects have 13, 48, 29, and 35 direct dependencies in their main modules, respectively. The developers of the project, go-admin-team/go-admin committed the lockfile at the early stages of the project but later decided to remove it from version control. The reason for removal is not documented in the public repository. The projects GoogleCloudPlatform/buildpacks and eolinker/apinto have go.sum files for some modules but do not contain one for the main module.
GoogleCloudPlatform/buildpacks also uses the Go vendor folder which allows committing dependencies alongside the project, making them accessible at anytime. Vendoring loads dependencies from the local vendor folder instead of resolving them from remote repositories, thus supporting air-gapped builds. Yet, in situations where the local cache cannot be trusted, the go.sum file would still be useful to ensure the integrity of the dependencies.
@@ 1. RQ3:
    What benefits do developers perceive in using lockfiles? Here, we study the benefits of lockfiles as perceived by developers and analyze how these perceptions differ across various package managers. Understanding the practical benefits of lockfiles helps clarify their role in secure and reliable software development.
@@ 2. RQ4:
    What challenges do developers face when managing lockfiles? With this RQ, we study the situations where developers struggle to maintain lockfiles, are unsatisfied with existing lockfile features, and find that lockfiles fall short of meeting their expectations. These findings are beneficial for package manager maintainers to improve their lockfile implementations for better usability and effectiveness.
4.2 Methodology for RQ3 and RQ4
To answer our 2 research questions on lockfile usage, we select a set of projects from each ecosystem, including some projects that publish their lockfile in the version control system and some projects that do not publish it. We then conduct semi-structured interviews with the top contributor of each project. In the following, we describe the methodology we follow to select the projects, contact developers, interview them and analyze their responses.
@@ 4.2 Methodology for RQ3 and RQ4
For each selected project, we gather publicly available contact information for the top contributor and send an interview invitation. If contact information is unavailable, or we do not receive a response within a week, we continue iterating through our randomly selected projects, until we reach our target of two or four participants per category, for a total of 14 responses. This approach allows us to select a diverse group of developers from various backgrounds and location, who have expertise in dependency management and who contribute to projects in various application domains.
@@
Interviews: We conduct 20 minute online interviews with 15 selected developers. We record the interviews upon consent from the participants. One interview (P8) was conducted in person at the participant’s request ; in this case, only written notes were taken. The participants were not provided any compensation for their participation.
@@
Interview response analysis: We use the Whisper speech recognition tool [18] to transcribe 15 interviews with available audio recordings. If audio recordings are not available, we refer to the written notes. Two authors carry out line-by-line coding for the transcribed interviews (or written notes). Consequently, the same two authors collaboratively create an affinity diagram, grouping codes into emerging themes inductively. The final affinity diagram consists of 20 sub-themes and 7 main themes related to general software development practices, lockfile benefits, developer expectations, and challenges.
Table 7: Demographics of developers interviewed about their usage of lockfiles. The project domains are selected based on GitHub project descriptions. The experience indicates the number of years of experience of participants in programming in any language.
Experience (Years) Lockfile Usage Project Stats Project Domain
@@
npm/pnpm P1 >{>}> 20 Yes 3k 110,892 320 Website development
@@
Cargo P6 2 - 5 Yes 1.9k 1338 50 Private social media
@@
Poetry Pipenv P10 5 - 10 Yes 834 512 27 Home Assistants
@@
4.3 Results for RQ3: What benefits do developers perceive in using lockfiles?
@@
4.3.1 Build determinism benefit
@@ 4.3.2 Integrity benefit
Identifying package tampering through integrity verification requires that releases remain immutable. In Go, when a package hosted on GitHub is modified after publication, such as through re-tagging or force-pushing, it can lead to checksum failures, as the version stored in the Go module proxy may differ from the modified release. A Go developer described this as a trade-off between the convenience of releasing and the security. (”it’s good on the one hand that you will always, like from a security perspective, make sure that you will always get the code that is basically stored on the proxy. And if you don’t have that, something is wrong and the checksum will give you an error. But then again, it can cause issues if you like want to re-release the same version. And yeah, it’s like convenience versus security”, P15)
4.3.3 Transparency benefit
@@ 4.3.5 Security benefit
4.4 Results for RQ4: What challenges do developers face when managing lockfiles?
@@
4.4.1 Indirect dependency locking challenge
@@
4.4.2 Slowing down dependency updates
@@
4.4.3 Human readability challenge
@@
When asked about the information they consider during code reviews of lockfiles, the majority of developers who commit lockfiles mention that they focus on the name, URL, version of the package, and the number of indirect dependencies it includes. A consistent observation we make is that none of the developers pay attention to the checksum of the dependencies in the lockfile. (I’ve never actually checked checksums, P7). Following the same reasoning, Go developers recommend the Go’s approach of having checksums in a separate file. (”To be honest, I really don’t care where the checksum is stored. It makes it clean to have it separated because I only have to put it into the Go mod file and then the checksum will be generated automatically in the sum file. So that’s like having a look at your Go mod file is much cleaner., P15 ”).
4.4.4 Cache challenge
@@ 4.4.5 Learning challenge
Findings from RQ4: Some developers find it challenging to manage lockfiles: readability challenges due to the verbosity of lockfiles, delays in dependency updates, or the challenge of enforcing the lockfile for a library. Overall, developers agree on the fact that lockfile challenges are better managed in Go than in JavaScript or Python.
5 Implications of Results
@@
5.1 Recommendations for package manager users
For developers who use package managers, we provide two main recommendations based on our analysis: selecting a package manager with a suitable lockfile feature, and following best practices for lockfile usage to get the most out of them.
[Recommendation #1] We recommend selecting package managers with human readable lockfiles that are enforced by default.
In the JavaScript ecosystem, we study two package managers: npm and pnpm. npm includes extensive metadata for each dependency, making its lockfile more verbose compared to pnpm. We observe that many developers struggle with maintaining and reviewing npm lockfiles. Therefore, we recommend pnpm for JavaScript developers who prefer manually monitoring lockfile changes.
The main difference between Poetry and Pipenv is in the lockfile-based resolution and enforcement. Poetry enforces the lockfile by default with the install action and throws an error if the lockfile needs to be updated whereas Pipenv silently updates the lockfile when a newer dependency version within the specified range in the dependency specification file is available. Therefore, we recommend Poetry over Pipenv for Python projects operating in untrusted environments, such as those relying on untrusted local caches, and requiring frequent dependency integrity checks.
[Recommendation #2] We recommend committing lockfiles for all projects.
Library developers face the same security risks and collaborative development challenges as application developers. As explained by library developers in subsection 4.4, the perceived cost of committing a lockfile is a false sense of security: they do not encounter in-range breakages themselves, while the library users still may. To mitigate such unexpected issues, library developers can run the library’s test suite both with and without enforcing locked versions in their CI/CD pipelines to detect any in-range breakages. Such an approach may allow them to commit a lockfile while still ensuring that the library remains free of in-range breakages. Another observation we made is that library developers often skip committing lockfiles because they believe determinism is not a primary concern for them. However, as discussed in subsection 4.3, lockfiles offer several other practical benefits such as enabling distributed integrity checks, which library developers can also take advantage of. Therefore, we recommend committing lockfiles for all projects.
5.2 Recommendations for package manager developers
Drawing inspiration from the best lockfile features, package manager developers can refine the lockfile functionality in existing package managers to improve the overall developer experience. Moreover, new package managers continue to emerge, introducing innovative features and addressing limitations in traditional ones. These new tools can adopt best practices around lockfiles and have the potential to influence a broad developer community. Based on our analysis, we suggest three recommendations to guide the future development of package managers.
[Recommendation #3] We recommend future package managers to prioritize developer experience.
@@ 5.2 Recommendations for package manager developers
[Recommendation #4] We recommend lockfiles to include only essential content.
When designing lockfiles, package managers should avoid duplicating information already present in the dependency specification file. They should also exclude data that is irrelevant to dependency locking. In npm, for example, including data such as licenses and funding information in the lockfile adds little to no perceived value while unnecessarily increasing code reviewing cost. At the other extreme, Gradle presents a serious lockfile anti-pattern by omitting essential details such as checksums. We recommend lockfiles to include the following mandatory details: resolved package versions, checksums, URLs for downloaded packages, and a clear distinction between direct and indirect dependencies. Any additional metadata should be kept as brief as possible, and ideally included only for reproducibility purposes, such as OS, engine, and language versions. We also encourage separating checksums into a separate file, similar to the approach of Go, to improve the human readability of lockfiles.
[Recommendation #5] We recommend package managers to generate lockfiles by default.
@@
6 Related work
To the best of our knowledge, our work is the first systematic analysis of lockfiles in the literature. Some studies occasionally mention them in the broader context of dependency management.
6.1 Dependency resolution and package managers
One of the key responsibilities of a package manager is resolving compatible versions for all the dependencies in a project. Zacchiroli et al.[19] review the usage of dependency resolving algorithms, including specialized constraint solving algorithms as well as ad-hoc dependency tree based algorithms. They conduct a census of dependency resolution mechanisms across 18 package managers, using version locking as one of the criteria for comparison. Pinckney et al.introduce a unified framework built on top of Max-SMT solvers to resolve dependencies more systematically, moving beyond ad-hoc algorithms [20].
Several studies focus on dependency conflict issues in different ecosystems. For example, [21] identifies factors that lead to conflicts, whereas [22] highlights issues related to inefficiency and excessive resource usage by the dependency resolution strategies in Python. [23] introduces a tool to generate crash stack traces for dependency conflicts in Java. Patra et al.[24] analyze dependency conflicts in JavaScript that arise from namespace collisions. [25] presents an in-depth study of conflicts between peer dependencies in npm. Similarly, in the Go ecosystem, Wang et al.[26] study conflicts caused by the coexistence of two library referencing modes: Go PATHs and Go modules.
Liu et al.[10] analyze vulnerability propagation within dependency trees by applying npm-specific dependency resolution rules, rather than relying on simple reachability analysis for indirect dependencies. They recommend managing dependencies with lockfiles. Along similar lines, He et al.[27] study the effect of dependency pinning in npm projects. They find that local pinning leads to more security vulnerabilities due to bloated and outdated dependencies. They also suggest that the risk of malicious package updates can be reduced when a small set of core dependencies pin their versions and keep them updated regularly.
@@ 6.1 Dependency resolution and package managers
6.2 Dependency management practices in open source software
Several empirical studies provide insights into the dependency management practices of open source software projects. Bogart et al.[28] discuss practices related to breaking changes across different package managers, and highlight the use of lockfiles to lock versions of indirect dependencies as a means to stabilize projects. Breaking updates have been well studied in the literature [29, 30, 6]. Venturini et al.[31] quantitatively evaluate the impact of breaking updates on dependent packages in the npm ecosystem. They also note the advantages of using lockfiles to track all resolved packages and avoid in-range breaking changes.
Schueller et al.[32] curate data from the Rust ecosystem over eight years, capturing developer activity, library dependencies, and usage trends. Another empirical analysis in Rust [33] reveals that 46% of packages adopted yanked releases and the proportion of yanked releases keeps increasing over the years. In Cargo, such yanked releases can be resolved only if a lockfile is present.
There is a body of work that analyzes security practices and threats associated with dependency management. Gu et al.[34] conduct a systematic analysis of security threats in package registries; Maven, PyPI, npm, Cargo, NuGet, and Go. Rahman et al.[35] propose a metric to evaluate vulnerabilities in dependencies based on the time it takes for a fix to be applied. Bos [36] reviews supply chain attacks and discusses a type of attack that originates from lockfile tamperings. One such tampering is adding a duplicate entry to the lockfile with a modified package and its new checksum. As most lockfiles are tediously long for thorough review, such tampering goes undetected leading to supply chain attacks. Vaidya et al.[37] characterize attacks in npm and PyPI ecosystems. Their findings suggest that with proper tooling and practices it is possible to mitigate these attacks to a great extent. The authors discuss the npm CLI lockfile npm-shrinkwrap that allows indirect dependency locking. They mention that npm-shrinkwrap can prevent future attacks from new packages through dependency locking. Kabir et al.[12] study security best practices in software development and consider committing lockfiles as one of the best practices. Their study focuses exclusively on the npm ecosystem, whereas our work spans multiple ecosystems and centers specifically on lockfiles.
@@ 6.2 Dependency management practices in open source software
6.3 Adoption of dependency management tools within development teams
Previous works have also used combined qualitative and quantitative methods to learn about developers’ adoption of dependency management tools. Schorlemmer et al.[38] study the adoption of software signing in four package registries; Maven, PyPI, DockerHub and Huggingface. Kalu et al.[39] conduct a follow-up interview study to understand software signing in practice. They identify that strict signature rules increase the quantity of signatures, and that registry policies have an impact on the developer decision to adopt proper signing.
Saroar et al.[40] survey 90 developers to understand how they use, choose, and troubleshoot GitHub Actions. Holtgrave et al.[41] analyze different attack types that can be performed by impersonating GitHub user accounts. Bifolco et al.[42] assess the accuracy of the GitHub dependency graph in Java and Python projects. Bifolco et al.use lockfiles as one source of truth to assess the GitHub dependency graph.
Dependency management bots such as dependabot, renovate and Greenkeeper automate the process of updating dependency specification files and lockfiles when new dependency versions are available. Alfadel et al.[43] and Mohayeji et al.[44] evaluate how developers respond to security updates suggested by dependabot. Rombaut et al.[45] study whether Greenkeeper reduces developer effort or introduces unnecessary workload. They mention lockfiles as a way to overcome in-range breaking changes, and also briefly mention about the challenge library developers face for sharing lockfiles. Our qualitative analysis confirms that developers do use bots to maintain dependency specification files and lockfiles up to date.